Four Students sit a typing test.
The probability that Anna passes the test is 1/2
The probablity that Boris passes the test is 4/9
Calulate the probabilty that both Anna and Boris pass the test. Give the answer in its simplest form?

Answer by stanbon(75887)

The probability that Anna passes the test is 1/2
The probablity that Boris passes the test is 4/9
Calulate the probabilty that both Anna and Boris pass the test. Give the answer in its simplest form?
-------------
The events are independent so you multiply to get:
P(both pass) = (1/2)(4/9) = 2/9
=================
Cheers,
Stan H.